# OASIS: Order-Augmented Strategy for Improved Code Search

### v3

@@ 1 Introduction
Code search tasks aim to retrieve the code snippet that best matches a given natural language (NL) query, thereby significantly enhancing developer productivity. A common approach is to leverage code embedding vectors to represent the semantics of the code for measuring its similarity to NL queries in code-NL matching tasks Nie et al. (2016); Husain et al. (2019); Shuai et al. (2020); Parvez et al. (2021); Di Grazia and Pradel (2023). Building on LLM advancements, code embeddings have significantly benefited recent code applications, including Retrieval-Augmented Generation (RAG) Asai et al. (2023); Gao et al. (2023) and code completion (ReAcc) Lu et al. (2022).
@@
However, most previous work relies on the difference between positive and negative pairs (i.e., major difference). This approach may result in superficial semantics due to the sparse nature of code context, where even a subtle change can lead to significant variations in functionality and meaning. To illustrate this insight, we present a sample from the CSN Python dataset in Figure 1. Focusing solely on major differences may lead to matching the query with code snippets 1 and 2, which only exhibit superficial similarity through word overlap.
To address this concern, we aim to research how to exploit subtle differences among negative pairs to improve code search. Prior work has primarily focused on intra-differences among negative pairs in two ways: some models have investigated “hard” negative pairs, which are difficult for models to differentiate Karpukhin et al. (2020); Gao et al. (2021a); Zhang et al. (2021), while others have adopted weighted optimization objectives to automatically balance learning from both hard and easy negatives Li et al. (2022b, 2023). In contrast, we propose OASIS (Order-Augmented Strategy for Improved code Search), which leverages order-based similarity labels to capture deeper semantic nuances. In this way, OASIS can better distinguish subtle differences among negative pairs through finer-grained comparisons for effective code search.
@@ 2 Related Work
OASIS is also related to broader code-related tasks. Inspired by the NLP paradigm, the typical practice involves representation learning through pre-training on a large code corpus, followed by fine-tuning for specific downstream tasks Feng et al. (2020); Wang et al. (2021). Here, code embedding is a crucial pre-training task, aiming to align code semantics with NL queries for code search. Following text embedding, they employ contrastive learning to explore code similarity Guo et al. (2022); Shi et al. (2023); Zhang et al. (2024). Some also incorporated code structures, such as data flow or Abstract Syntax Trees (AST), as auxiliary objectives Guo et al. (2020, 2022).
@@ 3.2 Similarity Refinement
Two methods are employed to extract candidate pairs requiring refinement. The first method filters candidate pairs based on the similarity scores, employing Gaussian Mixture Model (GMM) to fit the distribution of similarity scores across all sample pairs. Here, the similarity scores exhibit a bimodal distribution, with one peak corresponding to positive pairs and another to negative pairs. The intersection of these two distributions is taken as the threshold value for delineating positive and negative sample pairs. The equation is as below:
@@
Finally, an LLM is used to determine, with the docstring as a query, whether the candidate code or the code of the positive pair better answers the query. If LLM determines the candidate code can also satisfy docstring, then similarity of candidate pair is adjusted positively with , which is the optimal value attained by grid search. This approach is inspired by the robust performance of LLMs in binary choice tasks. Directly scoring pairs with an LLM typically results in most negative pairs’ scores clustering near zero, which makes it challenging to provide high-quality similarity adjustment.
@@ 3.3 Training Process of OASIS
Ultimately, two loss functions above are combined to form the overall optimization objective, with and serving as hyper-parameters to balance these objectives. The exact values of all hyper-parameters will be specified in the appendix.
@@ 4 Experimental Setup
In this section, we will provide a detailed elaboration of the experimental setup, including datasets, baselines, and evaluation metrics. An extensive training setting is available in the appendix.
@@ 5.3 Further Analysis
To provide a more detailed evaluation of OASIS’s performance within the test set, we extracted a hard subset from the CodeSearchNet dataset, specifically from the Python subset, comprising samples where all three models under study exhibited poor performance. These poorly performing samples are defined as instances where, given a query, the target code snippet does not rank first among the retrieved candidate snippets. OASIS utilizes order-augmented data, wherein sample pairs with different similarity labels are separately calculated for distance in comparison to positive sample pairs. Order-augmented approach enables the model to better discern which components contribute to the functionality required by the query. Consequently, the number of poorly performing samples in OASIS is relatively low, and in these samples, the rank of target code in retrieved candidates is generally higher. In the python subset, OASIS’s Mean Reciprocal Rank (MRR) significantly surpasses that of CodeSage-Large and Text-Embedding-3-Large, with the improvements of 5.46% and 5.35% as shown in Table 5. Other languages’ results can be found in the appendix.
@@ 6 Conclusion
Appendix A Training Settings
@@
For training model, the training of OASIS utilized Qwen/Qwen2.5-Coder-1.5B Team (2024) as the backbone model. For training data, the statistical details of the training dataset are presented in Table 2, which displays the quantity and proportion of data in different languages within the training dataset after processing through OASIS. Training and evaluation scripts are available at https://anonymous.4open.science/r/OASIS-E3EE.
@@
Appendix B Evaluation Dataset
@@
Appendix C Detailed Experimental Result
@@
Appendix D More Cases
Figures 5 and 6 illustrate two additional examples of similarity refinement. Figure 5 presents another example where the threshold filtering strategy was employed for similarity refinement. The docstring describes the functionality of the reference code as setting up a configuration and then performing assertions on the fields within this configuration dictionary. The candidate code also completes the configuration setup, as highlighted by the yellow rectangles in the figure, which mark sections of identical functionality. Although the candidate code tests different fields, it also includes assertion, thereby fulfilling the requirements described in the docstring. Consequently, the similarity score was increased from to .
@@ Appendix D More Cases
Figure 6 presents a case that triggered the AST filtering strategy. The reference code is derived from a solution to the programming challenge of Advent of Code 2020, Day 10, which aims to identify all possible adapter arrangements. The candidate code, predominantly using single-letter variable names, loses semantic information, resulting in a low initial similarity score of . However, the structural similarity between the candidate and reference codes, due to low AST edit distance, triggered the selection strategy. Upon confirmation through the LMM, the candidate code was an alternative implementation for the challenge, so the similarity was increased from to .

### v4

@@ Abstract
Zuchen Gao 1,2, Zizheng Zhan 3, Xianming Li 1, Erxin Yu 1
Ziqi Zhan, Haotian Zhang 3, Bin Chen 3, Yuqun Zhang 2, Jing Li 1††thanks: Jing Li is the corresponding author.
1 The Hong Kong Polytechnic University
2
Southern University of Science and Technology, 3
Kuaishou Technology.
{zuchen.gao, xianming.li, erxin.yu}@connect.polyu.hk
jing-amelia.li@polyu.edu.hk zhangyq@sustech.edu.cn
zhanziqi934@gmail.com, {zhanzizheng, zhanghaotian, chenbin}@kuaishou.com
https://huggingface.co/Kwaipilot/OASIS-code-embedding-1.5B
@@ 1 Introduction
Code search tasks aim to retrieve the code snippet that best matches a given natural language (NL) query, thereby significantly enhancing developer productivity. A common approach is to leverage code embedding vectors to represent the semantics of the code for measuring its similarity to NL queries in code-NL matching tasks Nie et al. (2016); Husain et al. (2019); Shuai et al. (2020); Parvez et al. (2021); Zeng et al. (2022); Di Grazia and Pradel (2023). Building on LLM advancements, code embeddings have significantly benefited recent code applications, including Retrieval-Augmented Generation (RAG) Asai et al. (2023); Gao et al. (2023) and code completion (ReAcc) Lu et al. (2022); Tan et al. (2024a) and repairing Xiang et al. (2024).
@@
However, most previous work relies on the difference between positive and negative pairs (i.e., major difference). This approach may result in superficial semantics due to the sparse nature of code context, where even a subtle change can lead to significant variations in functionality and meaning. To illustrate this insight, we present a sample from the CSN Python dataset in Figure 1. Focusing solely on major differences may lead to matching the query with code snippets 1 and 2, which only exhibit superficial similarity through word overlap. Detailed elaboration can be found in appendix A.
To address this concern, we aim to research how to exploit subtle differences among negative pairs to improve code search. Prior work has primarily focused on intra-differences among negative pairs in two ways: some models have investigated “hard” negative pairs, which are difficult for models to differentiate Karpukhin et al. (2020); Gao et al. (2021a); Zhang et al. (2021), while others have adopted weighted optimization objectives to automatically balance learning from both hard and easy negatives Li et al. (2022b, 2023); Zhuang et al. (2024). Details of existing methods for hard negative samples can be found in appendix B. In contrast, we propose OASIS (Order-Augmented Strategy for Improved code Search), which leverages order-based similarity labels to capture deeper semantic nuances. In this way, OASIS can better distinguish subtle differences among negative pairs through finer-grained comparisons for effective code search.
@@ 2 Related Work
OASIS is also related to broader code-related tasks. Inspired by the NLP paradigm, the typical practice involves representation learning through pre-training on a large code corpus, followed by fine-tuning for specific downstream tasks Feng et al. (2020); Tan et al. (2024b). Here, code embedding is a crucial pre-training task, aiming to align code semantics with NL queries for code search. Following text embedding, they employ contrastive learning to explore code similarity Guo et al. (2022); Shi et al. (2023); Zhang et al. (2024). Some also incorporated code structures, such as data flow or Abstract Syntax Trees (AST), as auxiliary objectives Guo et al. (2020, 2022).
@@ 3.2 Similarity Refinement
Two methods are employed to extract candidate pairs requiring refinement. The first method filters candidate pairs based on the similarity scores, employing Gaussian Mixture Model (GMM) to fit the distribution of similarity scores across all sample pairs. Here, the similarity scores exhibit a bimodal distribution, with one peak corresponding to positive pairs and the other to negative pairs. The intersection of these two distributions is taken as the threshold value for delineating positive and negative sample pairs. The equation is as below:
@@
Finally, an LLM is used to determine, with the docstring as a query, whether the candidate code or the code of the positive pair better answers the query. If LLM determines the candidate code can also satisfy docstring, then similarity of candidate pair is adjusted positively with , which is the optimal value attained by grid search. This approach is inspired by the robust performance of LLMs in binary choice tasks. Directly scoring pairs with an LLM typically results in most negative pairs’ scores clustering near zero, which makes it challenging to provide high-quality similarity adjustment. Elaboration with example is presented in appendix C.
@@ 3.3 Training Process of OASIS
Ultimately, two loss functions above are combined to form the overall optimization objective, with and serving as hyper-parameters to balance these objectives. The exact values of all hyper-parameters are specified in appendix E.
@@ 4 Experimental Setup
In this section, we will provide a detailed elaboration of the experimental setup, including datasets, baselines, and evaluation metrics. An extensive training setting is available in the appendix F.
@@ 5.2 Ablation Study
Thirdly, Table 4 also proved that our performance improvements are not solely attributed to LLM-generated docstrings. Our model achieves an NL2Code task performance of 65.49 when using only vanilla InfoNCE loss and the LLM-generated dataset (baseline loss with llm-generated docstring), which is even slightly below the average performance of CodeSage-Large (65.96). Notably, CodeSage’s dataset is derived from docstrings extracted directly from the original code using AST. The slightly lower performance could be attributed to the lack of weight term for negatives in the InfoNCE loss, potentially making the model susceptible to false negative pairs. When incorporating order-based loss on the same dataset, the model achieves a 1.8% improvement in MRR. Furthermore, the two distinct similarity refinement strategies each contributes additional improvements to the model’s overall 4.26% improvement, which proves the effectiveness of our method does not stem from high-quality LLM-generated docstrings.
Finally, as our model is slightly bigger, we conducted experiments on earlier and smaller-scale models to mitigate the influence of the model’s inherent capabilities. The results demonstrate that the performance improvement does not stem from the inherent strength of the backbone model itself but rather from the generalizability of the method. Appendix D presents a detailed analysis.
@@ 5.3 Further Analysis
To provide a more detailed evaluation of OASIS’s performance within the test set, we extracted a hard subset from the CodeSearchNet dataset, specifically from the Python subset, comprising samples where all three models under study exhibited poor performance. These poorly performing samples are defined as instances where, given a query, the target code snippet does not rank first among the retrieved candidate snippets. OASIS utilizes order-augmented data, wherein sample pairs with different similarity labels are separately calculated for distance in comparison to positive sample pairs. Order-augmented approach enables the model to better discern which components contribute to the functionality required by the query. Consequently, the number of poorly performing samples in OASIS is relatively low, and in these samples, the rank of target code in retrieved candidates is generally higher. In the python subset, OASIS’s Mean Reciprocal Rank (MRR) significantly surpasses that of CodeSage-Large and Text-Embedding-3-Large, with the improvements of 5.46% and 5.35% as shown in Table 5. Other languages’ results can be found in the appendix G.
@@ 6 Conclusion
Acknowledgments
This work is partially supported by the National Natural Science Foundation of China (Grant No. 62372220). Zuchen Gao, Xianming Li, Erxin Yu and Jing Li are substantially supported by Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25200821), the Innovation and Technology Fund (Project No. PRP/047/22FX), and PolyU Internal Fund from RC-DSAI (Project No. 1-CE1E). This work is also partially sponsored by CCF-Kuaishou Large Model Explorer Fund (NO. CCF-KuaiShou 2024013).
Appendix A Elaboration of Motivation Case
The Figure 1 demonstrates the failure to retrieve the correct result, the inability to effectively retrieve relevant results can be attributed to primarily focusing on superficial major features, such as keyword overlaps or approximate semantic matching. For instance, the code snippet ranked first was prioritized due to direct overlaps of keyword “parameter”, while the second-ranked snippet gained precedence by including more keyword “layer” (7, compared to 4 in the third-ranked example). In the top-2 cases, it appears that parameter names were matched to the query to some extent. However, subtle differences that are critical for semantic alignment were overlooked. Specifically, in rank 1, the snippet actually returns the column count or cached output rather than the parameter itself, and rank 2 also returns an output rather than the parameter. By contrast, the target code correctly returns the actual parameter, highlighting the inadequacy of existing methods in capturing these nuanced distinctions.
Appendix B Analysis of Existing Method for Hard Negative Samples
The methods of leveraging hard negative samples for training can be categorized into two approaches.
The first approach Karpukhin et al. (2020); Gao et al. (2021a); Zhang et al. (2021) involves selecting hard negative pairs by identifying sequences with high similarity to the current sequence based on metrics such as cosine similarity or BM25, while ensuring these sequences are actually dissimilar. These hard negative pairs are then combined with the original positive pairs and randomly selected negatives to construct the training dataset. During training, the loss function does not apply any special treatment to the hard negative samples. The model may be confused and struggle to identify the nuanced differences that decide whether it is the target code.
The second approach Li et al. (2022b, 2023); Zhuang et al. (2024) does not explicitly construct hard negative pairs. Instead, it introduces a weighting mechanism into the InfoNCE loss, where the model assigns lower weights to hard negatives. Essentially, this approach ignores the high similarity of hard negatives and treats them as regular negative samples. Consequently, the observed performance improvement may stem from mitigating the adverse impact of false negatives on the model.
Appendix C A Detailed Elaboration with Examples for Main Method
The case presented in the introduction demonstrates that existing methods struggle to capture subtle differences, particularly for code snippets that appear to match the query but are, in fact, dissimilar. These code snippets along with the query comply with the definition of hard-negative pairs. Traditional approaches for incorporating hard negative pairs suffer from their respective limitations. To address this, in our method, after Step 3.1, we obtain triplet data in the form of , where each natural language (NL) or code query is associated with multiple code snippets. For a given NL query, the dataset contains multiple pairs with different code with varying similarity scores, such as:
The similarity of 0.6 is annotated by embedding model, it is manually set to 1.0 during training as it is a positive pair.
While and are not the target code, their similarity scores with remain relatively high compared to unrelated samples. This indicates that and constitute hard negative pairs. Unlike traditional methods, our approach has two key features:
Preserving Relative Similarity through Order Loss: Instead of treating hard negative pairs as purely negative samples, the order-based loss leverages the relative similarity among these pairs. It trains the model to preserve the order relationship of similarity scores, capturing the subtle distinctions between them.
Progressive Learning with Hard Negatives: Each NL or code query forms multiple hard negative pairs with varying degrees of similarity, creating a sequence of progressively harder negative samples closer to the positive target. For instance, the model may learn that the dissimilarity between and the target code arises from two minor differences, whereas the dissimilarity between and the target code stems from only one small detail. This progressive learning enables the model to focus on the nuanced variations that distinguish hard negatives from the target code, rather than fully regarding them as irrelevant negatives. By doing so, the proposed framework effectively captures these fine-grained differences and enhances the model’s ability to distinguish between similar yet non-identical code snippets.
C.1 Concrete details of two methods (filtering strategies)
To fully utilizing hard negative pairs with order-based loss, it is essential to ensure that the similarity labels provided have a high degree of accuracy. However, the labels generated by other embedding models are often suboptimal and require more fine-grained adjustments. To address this, we propose two distinct strategies for selecting pairs that require refinement. Using the earlier example of:
The two strategies are as follows:
Threshold-Based Suspicious Candidate Filtering: If, for a given NL-query (e.g., ), any negative pair with exhibits a similarity score that exceeds the original annotated similarity of the positive pair (e.g., ) or crosses the threshold where the distribution of positive and negative samples overlaps (e.g., ), such pairs are chosen as suspicious candidate pairs. Because these pairs exhibit a higher similarity compared to the positive pair or compared to the statistical boundary of average positive and negative pairs. This indicates that these negative pairs may require further adjustment.
Structural Similarity-Based Filtering: Even if a negative sample has a low similarity score (e.g., with ), it is also marked as a suspicious candidate pair if its Abstract Syntax Tree (AST) exhibits a certain degree of structural similarity to the AST of the target code (e.g., ). This suggests that structurally similar negative samples may warrant additional adjustment.
The flagged suspicious candidate pairs (e.g., ) are then input to LLM alongside the positive pair (e.g., ). The model is tasked with determining whether the negative samples (e.g., ) better address the NL-query (e.g., ) in some way. If the model confirms this hypothesis, the original similarity score is adjusted by applying a small positive offset proportional to a scaling factor , such that:
This adjustment process refines the similarity scores, leading to improved overall performance. The purpose of both strategies is to identify and flag potentially inaccurate suspicious candidate pairs for further processing and refinement.
Model CoSQA AdvTest CSN Python Java JS PHP Go Ruby AVG Qwen2.5-Coder-1.5B (w/o training) OASIS (codebert-mlm) (125M) CodeSage-small (130M)
Appendix D Smaller Backbone Model
The performance margin may source from the capacity and scale of the backbone model, and Qwen2.5-Coder-1.5B is indeed a powerful backbone model and it is slightly bigger than CodeSage-Large (1.3B). To guarantee the generalization of our method, we conducted experiments when training on small-scale and earlier non-decoder backbone model CodeBERT Feng et al. (2020).
Table 6 shows the results of NL2Code tasks when using codebert-base-mlm as backbone model. We use the hidden states of cls token as the embedding. It clearly shows that the backbone model Qwen2.5-Coder-1.5B w/o training lacks the capability to perform retrieval tasks. When training with backbone model of codebert, OASIS can still outperform CodeSage-small in the same scale. With a margin of 2% on MRR on CSN, 6.8% on CoSQA, and 1.21% on AdvTest, which proves the effectiveness of our method on different scales of models.
Appendix E Training Settings
@@
For training model, the training of OASIS utilized Qwen/Qwen2.5-Coder-1.5B Team (2024) as the backbone model. For training data, the statistical details of the training dataset are presented in Table 2, which displays the quantity and proportion of data in different languages within the training dataset after processing through OASIS. Evaluation scripts are available at https://github.com/Zuchen-Gao/OASIS.
@@
Model CoSQA AdvTest CSN Avg Python Java JS PHP Go Ruby OASIS () OASIS () OASIS ()
Num | CoSQA | AdvTest | CSN | |||||
Python | Python | Python | Java | JS | PHP | Go | Ruby | |
Query | ||||||||
Candidate |
Num | Python | Java | JS | TS | C# | C | Ruby | PHP | GO |
Query | |||||||||
Candidate |
This result explains why when adjusting similarity judgments, we retain the original positive/negative role to avoid performance degradation caused by fluctuations in LLM outputs. This decision ensures that the labeling quality of the dataset is not entirely dictated by the LLM. As demonstrated in Table 7, when the is increased to 0.2—allowing for greater LLM intervention—performance actually deteriorates.
Appendix F Evaluation Dataset
@@
Appendix G Detailed Experimental Result
@@
Appendix H Choice of Embedding Model
The accuracy of code similarity annotations significantly affects the final performance of the model. Different embedding models often exhibit distinct similarity distributions. For example, Text-Embedding-3-Large tends to assign similarity scores around 0.5 for positive pairs and approximately 0.05-0.3 for negative pairs. In contrast, gte-Qwen1.5-7B-instruct assigns similarities around 0.4 for positive pairs and 0.2 for negative pairs, resulting in a narrower similarity range. However, while the absolute similarity scores vary across different embedding models, the order of similarity scores within a sample group remain the same. For instance, consider a positive pair and three corresponding negative pairs:
Here, the similarity score 0.6 for the positive pair is manually adjusted to 1.0 during training. The range of similarity scores (max-min) is 0.5. When switching to another embedding model, the similarity scores for the same group may shrink to a narrower range, such as:
In this case, the range is reduced to 0.3. Yet, the relative order of similarity scores within the group remains consistent. Since the loss function depends on the relative ranking of similarity scores rather than their absolute values, using different embedding models may lead to changes in score variance but does not affect the loss computation or the final training outcomes. To empirically verify that the similarity rankings annotated by different embedding models are approximately consistent, we evaluated two alternative models by re-annotating similarity scores for the same dataset and measuring the rank correlation. Specifically, for each NL, the similarity scores of its corresponding code pairs (e.g., , , etc.) were re-annotated using the alternative models, and the ranking consistency of a group was assessed using the nDCG metric. The evaluation was conducted on a randomly selected subset of 500,000 samples, containing approximately 90,000 unique NL-code groups. The average nDCG scores for the two alternative models are as follows:
gte-Qwen2-1.5B-instruct:
e5-mistral-7b-instruct:
These results demonstrate that when using different embedding models for similarity annotation, the relative ranking of similarity scores within the group of pairs with the same NL remain the same. As a result, the training process and final loss remain consistent regardless of the embedding model used for annotation.
Appendix I More Details
Computational Cost
For similarity annotation cost, due to separate annotation process, only a rough estimate of the time can be provided. Overall, it took approximately 7 days to complete the similarity annotation for the entire 53M dataset on 4 nodes (100 multi-process each). After the similarity annotations and adjustments were finalized, there was no additional computational overhead during the training phase. This is because order-based loss only relies on the hidden states from the final layer of the model for computation, ensuring that training efficiency remains unaffected.
Details about similarity refinement
For similarity refinement, the threshold-based strategy required approximately 84 hours (3.5 days) to complete, whereas the AST-based strategy took around 60 hours (2.5 days). The refinement process was conducted on a single node with 100 processes. The final training dataset comprises 53,825,502 samples, of which 647,521 samples (1.2%) were refined using the threshold strategy, 115,689 samples (0.2%) were refined using the AST strategy, and the 98.6% of the samples remained unchanged.
Appendix J More Cases
@@ Appendix I More Details
Figures 5 and 6 illustrate two additional examples of similarity refinement. Figure 5 presents a case that triggered the AST filtering strategy. The reference code is derived from a solution to the programming challenge of Advent of Code 2020, Day 10, which aims to identify all possible adapter arrangements. The candidate code, predominantly using single-letter variable names, loses semantic information, resulting in a low initial similarity score of . However, the structural similarity between the candidate and reference codes, due to low AST edit distance, triggered the selection strategy. Upon confirmation through the LMM, the candidate code was an alternative implementation for the challenge, so the similarity was increased from to .
Figure 6 presents another example where the threshold filtering strategy was employed for similarity refinement. The docstring describes the functionality of the reference code as setting up a configuration and then performing assertions on the fields within this configuration dictionary. The candidate code also completes the configuration setup, as highlighted by the yellow rectangles in the figure, which mark sections of identical functionality. Although the candidate code tests different fields, it also includes assertion, thereby fulfilling the requirements described in the docstring. Consequently, the similarity score was increased from to .